### v1

@@ 2.4 ZERO TRUST PRINCIPLES AND EVOLUTION
Traditional perimeter-based security models, which assume trust within a protected network, are no longer sufficient for modern, interconnected systems like autonomous vehicles. To address this, the National Institute of Standards and Technology (NIST) formalized the concept of Zero Trust Architecture (ZTA) in its Special Publication 800-207, establishing the guiding principle of "never trust, always verify" [10]. Under this model, every access request, regardless of origin, must undergo continuous authentication and authorization. This paradigm shift has been adopted by critical sectors, including the U.S. Department of Defense [54], and is particularly relevant for AVs, which operate in highly dynamic environments where static trust assumptions are impractical [55, 9, 5].
@@
Despite its advantages, ZTA faces implementation challenges when applied to vehicular and IoT systems. Annabi et al. [56] provide a comprehensive review of ZTA in connected vehicles, identifying open challenges in trust establishment, latency, and integration with distributed technologies. He et al. survey the state of ZTA, noting barriers such as system complexity, skill gaps, and interoperability issues, while also identifying blockchain as a potential enabler [12]. Li et al. extend this perspective to the Future Internet of Things (FIoT), arguing that ZTA can mitigate risks in massive device deployments combined with continuous verification and blockchain-assisted authentication [57]. The core principles of continuous verification and least privilege, however, align directly with the needs of AV ecosystems, making ZTA a foundational model that requires integration with decentralized trust anchors to overcome its inherent limitations [5].
@@ 2.5 ZERO TRUST AND BLOCKCHAIN INTEGRATION
While ZTA provides a strong conceptual foundation through continuous verification and least privilege, its centralized decision-making components face challenges in highly dynamic environments. Blockchain emerges as a natural complement to ZTA, addressing weaknesses related to central trust, credential revocation, and tamper-proof logging [58, 59, 60]. Its decentralized and immutable nature enables auditable access decisions and distributed identity verification, eliminating single points of failure. This synergy is particularly potent in the context of decentralized identity (DID) and verifiable credentials (VCs). DID frameworks allow entities to establish and prove identity without relying on centralized certificate authorities, while VCs enable selective disclosure and privacy-preserving proofs of authorization. These approaches align naturally with ZTA’s principle of continuous verification, where every request is evaluated independently of network location or prior trust [61, 62, 63].
@@
The integration of these paradigms has been explored across IoT and edge computing contexts, which share similar constraints with vehicular networks. For example, Dhar and Bose [58] highlight blockchain’s role in securing IoT devices under ZTA, and Li et al. [64] propose a blockchain-based zero-trust scheme for edge environments, addressing lightweight identity verification. n next-generation networks, Maksymyuk et al. [65] envision a blockchain-empowered framework for decentralized management in 6G, while Hassan et al. [66] apply blockchain-enabled Zero Trust identity management to smart cities and IoT networks. These frameworks provide a blueprint for applying blockchain-enhanced ZT in distributed and resource-constrained ecosystems. In the vehicular domain, numerous frameworks have been proposed for blockchain-based identity management, authentication, and decentralized trust in IoV and D-V2X [67, 68, 69, 70, 71, 72, 73]. For instance, George et al. proposed a secure identity management framework for VANETs, while Noh et al. developed a distributed message authentication scheme.
2.6 BLOCKCHAIN IN VEHICULAR AND ITS SECURITY
@@
Blockchain has emerged as a promising enabler for Intelligent Transportation Systems (ITS), offering decentralized trust, tamper-resistant logging, and auditable data exchange. Early visionary work by Yuan and Wang [74] highlighted the potential of blockchain to address trust and coordination challenges in ITS. Subsequent surveys and frameworks confirmed these prospects, exploring applications ranging from secure communication to resource management [75, 76, 77, 78]. Collectively, these studies frame blockchain as a foundational technology for achieving decentralized trust in connected transportation ecosystems.
@@
In the vehicular networking domain, blockchain has been proposed as a decentralized trust anchor for secure message exchange and system integrity. Distributed trust management schemes demonstrate how blockchain can mitigate single points of failure and enhance transparency in VANET operations [79]. Beyond fundamental networking, blockchain has supported a wide range of cooperative vehicular applications. For instance, Li et al. [80] proposed a blockchain-based cooperative perception framework in the Internet of Vehicles (IoV), enabling vehicles to securely share sensory information. Song et al. [81] developed a blockchain-enabled cooperative positioning system that integrates deep learning to enhance location accuracy. In the forensic domain, Cebe et al. [82] introduced Block4Forensic, a lightweight blockchain framework for accident investigation, while Zhu et al. [83] presented a blockchain-based accident forensics system for smart connected vehicles.
@@
Authentication and access control have been a central focus of blockchain-based vehicular security. Tan and Chung [69] proposed a secure blockchain-assisted key management scheme for VANETs, while Noh et al. [68] designed a distributed blockchain-based message authentication protocol. Adja et al. [84] advanced certificate revocation and status verification using blockchain, and Das et al. [70] developed a blockchain vehicle identity management framework for ITS.
Figure 3: Zero-Trust Decentralized Blockchain-based Communication Framework for Autonomous Vehicles.
@@
Despite these advances, scalability and performance remain persistent challenges. Blockchain consensus mechanisms can introduce significant computational and communication overhead, constraining their use in highly mobile vehicular environments. Dwivedi et al. [85] proposed batch authentication to reduce verification overhead, while Adja et al. [84] and Liu et al. [60] proposed lightweight revocation and information sharing mechanisms to enhance efficiency. These approaches illustrate efforts to reconcile blockchain’s benefits with the stringent performance requirements of vehicular networks.
@@
Despite extensive progress in identity management, Zero Trust, and blockchain-enabled ITS security, no integrated solution currently exists that combines Zero Trust principles with blockchain-based decentralized identity management tailored to autonomous vehicles. Existing solutions fail to simultaneously ensure scalability, timeliness of revocation, privacy-preserving authentication, and continuous verification in dynamic vehicular networks with a minimal overhead for real-time safety-critical applications such as platooning and cooperative perception. Our work addresses this gap by proposing a Zero Trust-based Decentralized Identity Management (D-IM) framework. By combining blockchain’s decentralized and tamper-resistant architecture with Zero Trust’s continuous verification and least-privilege principles, the proposed framework delivers a resilient, scalable, and privacy-aware identity management solution tailored to the stringent requirements of autonomous vehicles.
@@
3 DECENTRALIZED IDENTITY MANAGEMENT (D-IM) SYSTEM DESIGN
Autonomous vehicles operating in Vehicle-to-Everything (V2X) environments, where vehicles and roadside units (RSUs) must authenticate each other in real time before exchanging data, must continuously exchange safety-critical information such as location, speed, and intent. Guaranteeing the authenticity of these messages requires robust identity management. Traditional PKI- and pseudonym-based systems face scalability, revocation, and centralization issues, as discussed earlier in Section 2. To address these challenges, we propose a decentralized identity management (D-IM) system that integrates blockchain as a distributed trust anchor with Zero Trust principles of continuous verification and least-privilege access. This system eliminates reliance on centralized authorities, ensures continuous verification, and leverages blockchain for immutable identity records. This section presents the overall system model, attacker model, the D-IM communication framework, and the blockchain-based identification and authorization protocol.
3.1 SYSTEM MODEL
The proposed D-IM system is designed for V2X communication environments consisting of three main components:
  • Participants (vehicles and RSUs): The primary participants that act as both data producers and consumers. Participants must authenticate with each other before exchanging any information.
  • Verification Nodes (RSUs): Act as verification and validation nodes. RSUs maintain blockchain connectivity, assist in identity verification, and participate in consensus processes.
  • Blockchain Network: A decentralized ledger that stores identity credentials (IDs, public keys, timestamps, location metadata, and cryptographic hashes). It serves as the immutable trust anchor for identity management.
Figure 3 presents the overall Zero Trust decentralized blockchain-based communication framework. Each vehicle and RSU runs a local copy of the framework code and communicates with the blockchain to achieve secure and verifiable V2X communications. The framework enforces Zero Trust by requiring identity verification for every interaction, eliminating reliance on any single trusted entity.
@@ 3.1 SYSTEM MODEL
11:   Both vehicles compute $K_{s}$ from $Q$ and $R$ via Diffie–Hellman
@@
3.2 ZERO TRUST DECENTRALIZED COMMUNICATION ARCHITECTURE
The Zero Trust decentralized communication framework requires every identity in the network is verified before any data exchange. Vehicles and RSUs interact with the blockchain to obtain authenticated identity information. RSUs, acting as verification nodes, participate in validating credentials and anomaly detection. Suspicious behaviors are flagged, and verification nodes may collaborate to reach consensus on whether a participant is misbehaving.
The overall communication framework (Figure 3) integrates Zero Trust principles into V2X security. Each vehicle and RSU interacts with the blockchain to: (i) register identities in a decentralized manner, (ii) continuously verify peer authenticity, (iii) analyze communication behavior to detect anomalies, and (iv) enforce dynamic security policies. Unlike perimeter-based security models, every access attempt is verified without assumption of trust, ensuring resilience against insider and outsider threats.
An overview of the system steps is shown in Figure 4. Authentication via blockchain precedes all information sharing, ensuring that only legitimate participants contribute to cooperative V2X operations. Smart contracts could be used to automatically penalize cheating behavior and enforce fairness, such extensions are left as future work.
3.3 BLOCKCHAIN-BASED IDENTIFICATION AND AUTHORIZATION PROTOCOL
The protocol operates as follows:
@@ 3.3 BLOCKCHAIN-BASED IDENTIFICATION AND AUTHORIZATION PROTOCOL
    Initialization: Each vehicle submits its ID, public key, timestamp, location, and the hash of its ID with its private key to the blockchain.
@@
    Request/Response: When Vehicle 1 requests to communicate with Vehicle 2, it queries the blockchain for Vehicle 2’s credentials. Similarly, Vehicle 2 queries the blockchain for Vehicle 1’s information.
@@
    Verification: Each vehicle cross-verifies the other’s public key and hash against blockchain records to ensure authenticity.
@@
    Session Key Establishment: Both vehicles exchange random values ($Q$, $R$) encrypted with the other’s public key, enabling them to compute a shared session key $K_{s}$ using the Diffie–Hellman method.

### v2

@@ 2.2 Zero Trust Principles and Evolution
Identity management is a cornerstone of secure V2X communications, ensuring that only authenticated entities can participate in the exchange of safety-critical messages. Formal frameworks have been established by standardization bodies like IEEE STD 1609.2 [7426684] and ETSI TS 103 097 [etsi:103097], which form the basis of Public Key Infrastructure ,PKI, based vehicular credential management systems. In this approach, vehicles obtain long-term certificates from a root authority and use short-term pseudonym certificates to preserve privacy and unlinkability [petit2014pseudonym]. However, this anonymity comes at a cost, as regulatory and forensic requirements often necessitate conditional traceability, creating a tradeoff between privacy and accountability [khodaei2015key, petit2014pseudonym, yoshizawa2023survey].
Scalability and privacy are also ongoing limitations. High pseudonym change rates strain storage and distribution systems, while insufficient changes increase linkability risks [yoshizawa2023survey]. Recent surveys emphasize that current PKI-based identity frameworks struggle to balance these needs [ying2024literature, yoshizawa2023survey, almarshoud2024security, farsimadan2025review, ying2024literature, kifor2024automotive]. These limitations highlight the need for more robust, decentralized approaches to identity management, motivating the exploration of blockchain and Zero Trust principles as complementary paradigms.
Traditional perimeter-based security models, which assume trust within a protected network, are no longer sufficient for modern, interconnected systems like autonomous vehicles. To address this, the National Institute of Standards and Technology (NIST) formalized the concept of Zero Trust Architecture (ZTA) in its Special Publication 800-207, establishing the guiding principle of "never trust, always verify" [nist800207]. Under this model, every access request, regardless of origin, must undergo continuous authentication and authorization. This paradigm shift has been adopted by critical sectors, including the U.S. Department of Defense [dod2022zt], and is particularly relevant for AVs, which operate in highly dynamic environments where static trust assumptions are impractical [wylde2021zero, anderson2023zero, yousseef2025autonomous].
@@
Despite its advantages, ZTA faces implementation challenges when applied to vehicular and IoT systems. Annabi et al. [annabi2025survey] provide a comprehensive review of ZTA in connected vehicles, identifying open challenges in trust establishment, latency, and integration with distributed technologies. He et al. survey the state of ZTA, noting barriers such as system complexity, skill gaps, and interoperability issues, while also identifying blockchain as a potential enabler [he2022survey]. Li et al. extend this perspective to the Future Internet of Things (FIoT), arguing that ZTA can mitigate risks in massive device deployments combined with continuous verification and blockchain-assisted authentication [li2022future]. The core principles of continuous verification and least privilege, however, align directly with the needs of AV ecosystems, making ZTA a foundational model that requires integration with decentralized trust anchors to overcome its inherent limitations [yousseef2025autonomous].
@@ 2.3 Zero Trust and Blockchain Integration
While ZTA provides a strong conceptual foundation through continuous verification and least privilege, its centralized decision-making components face challenges in highly dynamic environments. Blockchain emerges as a natural complement to ZTA, addressing weaknesses related to central trust, credential revocation, and tamper-proof logging [dhar2021securing, alevizos2022augmenting, liu2022blockchain]. Its decentralized and immutable nature enables auditable access decisions and distributed identity verification, eliminating single points of failure. This synergy is particularly potent in the context of decentralized identity (DID) and verifiable credentials (VCs). DID frameworks allow entities to establish and prove identity without relying on centralized certificate authorities, while VCs enable selective disclosure and privacy-preserving proofs of authorization. These approaches align naturally with ZTA’s principle of continuous verification, where every request is evaluated independently of network location or prior trust [vcuvcko2021decentralized, dib2020decentralized, stockburger2021blockchain].
@@
The integration of these paradigms has been explored across IoT and edge computing contexts, which share similar constraints with vehicular networks. For example, Dhar and Bose [dhar2021securing] highlight blockchain’s role in securing IoT devices under ZTA, and Li et al. [li2022zero] propose a blockchain-based zero-trust scheme for edge environments, addressing lightweight identity verification. n next-generation networks, Maksymyuk et al. [maksymyuk2020blockchain] envision a blockchain-empowered framework for decentralized management in 6G, while Hassan et al. [hassan2023blockchain] apply blockchain-enabled Zero Trust identity management to smart cities and IoT networks. These frameworks provide a blueprint for applying blockchain-enhanced ZT in distributed and resource-constrained ecosystems. In the vehicular domain, numerous frameworks have been proposed for blockchain-based identity management, authentication, and decentralized trust in Internet of Vehicles (IoV) and V2X [george2020secure, noh2020distributed, tan2019secure, das2023secure, theodouli2020towards, agudo2020blockchain, van2017blackchain]. For instance, George et al. proposed a secure identity management framework for VANETs, while Noh et al. developed a distributed message authentication scheme.
Despite these advances, performance and scalability remain significant concerns. Blockchain protocols often introduce overhead that is ill-suited to the high mobility and low-latency demands of vehicular networks. This is due to the computational cost of cryptographic operations, the communication overhead of broadcasting transactions, and the latency of consensus protocols. Researchers have therefore focused on lightweight mechanisms for revocation and verification to balance security with real-time performance. Adja et al. [adja2021blockchain] advanced certificate revocation using blockchain, while Liu et al. [liu2022blockchain] proposed a decentralized, fair, and authenticated information-sharing scheme for the Internet of Things. These efforts, along with works on decentralized network management in 6G [maksymyuk2020blockchain], illustrate the ongoing challenge of reconciling blockchain’s benefits with the stringent performance requirements of vehicular networks. In summary, while ZTA provides the principles and blockchain offers decentralized trust, their combined use requires optimization to address the scalability and latency challenges of current vehicular networks, motivating the framework presented in this paper.
2.4 Blockchain in Vehicular Networks and its Security
@@
Blockchain has emerged as a promising enabler for Intelligent Transportation Systems (ITS), offering decentralized trust, tamper-resistant logging, and auditable data exchange. Early visionary work by Yuan and Wang [yuan2016towards] highlighted the potential of blockchain to address trust and coordination challenges in ITS. Subsequent surveys and frameworks confirmed these prospects, exploring applications ranging from secure communication to resource management [cocirlea2020blockchain, mollah2020blockchain, jabbar2022blockchain, das2023blockchain]. Collectively, these studies frame blockchain as a foundational technology for achieving decentralized trust in connected transportation ecosystems.
@@
In the vehicular networking domain, blockchain has been proposed as a decentralized trust anchor for secure message exchange and system integrity. Distributed trust management schemes demonstrate how blockchain can mitigate single points of failure and enhance transparency in VANET operations [inedjaren2021blockchain]. Beyond fundamental networking, blockchain has supported a wide range of cooperative vehicular applications. For instance, Li et al. [li2021blockchain] proposed a blockchain-based cooperative perception framework in the IoV, enabling vehicles to securely share sensory information. Song et al. [song2020blockchain] developed a blockchain-enabled cooperative positioning system that integrates deep learning to enhance location accuracy. In the forensic domain, Cebe et al. [cebe2018block4forensic] introduced Block4Forensic, a lightweight blockchain framework for accident investigation, while Zhu et al. [zhu2024blockchain] presented a blockchain-based accident forensics system for smart connected vehicles.
@@
Authentication and access control have been a central focus of blockchain-based vehicular security. Tan and Chung [tan2019secure] proposed a secure blockchain-assisted key management scheme for VANETs, while Noh et al. [noh2020distributed] designed a distributed blockchain-based message authentication protocol. Adja et al. [adja2021blockchain] advanced certificate revocation and status verification using blockchain, and Das et al. [das2023secure] developed a blockchain vehicle identity management framework for ITS.
Blockchain has also been leveraged to ensure data integrity and trustworthy event reporting. Dwivedi et al. [dwivedi2021blockchain] proposed an IPFS-integrated blockchain protocol for secure event storage and sharing in VANETs, while Vangala et al. [vangala2020blockchain] introduced the BCAS-VADN system to authenticate accident detection and notification reports on the blockchain. Resource sharing and content distribution also benefit from blockchain. Miglani and Kumar [miglani2023blockchain] developed a blockchain-based framework for content-centric vehicle-to-grid (V2G) networks, ensuring secure and efficient content dissemination. Theodouli et al. [theodouli2020towards] extended blockchain to identity and trust management in IoV ecosystems, while Stockburger et al. [stockburger2021blockchain] explored decentralized identity for public transportation.
@@
Scalability and performance remain persistent challenges in blockchain-enabled vehicular security systems, as consensus mechanisms can introduce significant computational and communication overhead that constrains their applicability in highly mobile V2X environments. To mitigate these limitations, Dwivedi et al. [dwivedi2023design] proposed batch authentication techniques to reduce verification overhead, while Adja et al. [adja2021blockchain] and Liu et al. [liu2022blockchain] introduced lightweight revocation and information-sharing mechanisms to improve efficiency. Collectively, these efforts highlight ongoing attempts to reconcile the security benefits of blockchain with the stringent performance requirements of vehicular networks.
@@
Although substantial progress has been made in identity management, Zero Trust architectures, and blockchain-enabled ITS security, an integrated solution that combines Zero Trust principles with blockchain-based decentralized identity management specifically tailored to autonomous vehicles is still lacking. Existing approaches do not simultaneously achieve scalability and privacy-preserving authentication under the tight latency constraints imposed by dynamic vehicular networks and real-time safety-critical applications such as platooning and cooperative perception.
3 DECENTRALIZED IDENTITY MANAGEMENT (D-IM) SYSTEM Model
Autonomous vehicles operating in Vehicle-to-Everything, V2X, environments, where vehicles and roadside units (RSUs) must authenticate each other in real time before exchanging data, must continuously exchange safety-critical information such as location, speed, and intent. Guaranteeing the authenticity of these messages requires robust identity management. Traditional PKI- and pseudonym-based systems face scalability, and centralization issues, as discussed earlier in Section 2. To address these challenges, we propose a decentralized identity management (D-IM) system that integrates blockchain as a distributed trust anchor with Zero Trust principles of continuous verification and least-privilege access. Our proposed system eliminates reliance on centralized authorities, ensures continuous verification, and leverages blockchain for immutable identity records. This section presents the overall system model, attacker model, the D-IM communication framework, and the blockchain-based identification and authorization protocol.
3.1 Design Goals
The design goals of our approach are as follows:
  • Mutual Authentication: In a zero trust environment, to ensure the identities of sharing parties, we verify their authenticity through blockchain technology after authenication of each contributing party. This ensures that only legitimate participants are allowed to share information.
  • Autonomy: The authentication and authorization process, consensus mechanism and information sharing process are governed by blockchain technology, eliminating the need for any third-party authority or centralized entity.
  • Anonymity: Participants may be concerned about their personal information (such as geographical positions and other sensitive or identifying details) being exposed during the information-sharing process. Therefore, blockchain is utilized to enable the sharing of real-time traffic information without revealing the participant’s true identity since a pseudo-identity is given to each vehicle after authentication success. Only the blockchain distributed ledger stores the real identity information.
  • Privacy Protection: The shared information should not compromise the privacy of the authenticated vehicles, i.e. disclosing their location, to unauthenticated vehicles. To ensure this, the information is always encrypted.
  • Non-Repudiation and Traceability of Origin: The IDs and keys used will be stored on the blockchain distributed ledger to trace the history of information sharing.
3.2 System Model
The proposed D-IM system is designed for V2X communication environments consisting of two main components:
  • Blockchain Network: A decentralized ledger that stores identity credentials (IDs, public keys, timestamps, location metadata, and cryptographic hashes). It serves as the immutable trust anchor for identity management. It consists of the Verification Peers (RSUs). These verification validation peers hold the blockchain ledger records, maintain blockchain connectivity, assist in identity verification, and participate in consensus processes.
  • Participants Vehicles: The primary participants that act as both data producers and consumers. Participants must be authenticated and verified before exchanging any information with each other or with the blockchain network. We are planning to utilize machine learning based intrusion detection system [abrar2024gps] and blockchain smart contracts to ensure fairness and penalize malicious behavior, revoke untrusted participants from the network and retract anomalous, false or misleading shared information as a future extension work based on this paper.
The proposed Zero Trust decentralized communication framework requires every identity in the network is verified before any data exchange. Vehicles and RSUs interact with the blockchain to obtain authenticated identity information. RSUs, acting as verification nodes, participate in validating credentials and anomaly detection. Suspicious behaviors are flagged, and verification nodes may collaborate to reach consensus on whether a participant is misbehaving.
3.3 Attacker Model
We assume an adversarial environment where attackers may attempt to compromise vehicular communications. Specific attack vectors include:
  1. 1.
    Identity and Authentication Attacks:
    • Masquerading/Impersonation Attack: occurs when an attacker compromises a node (such as an Electronic Control Unit or an entire vehicle) and uses its legitimate identity to send malicious messages, making them appear genuine to other systems. The attacker then uses the gained access for eavesdropping or sending false sensors data, i.e. position/velocity data falsification.
    • Sybil Attacks: are when one attacker creates multiple fake identities (Sybil nodes) to broadcast false data (like fake accidents, traffic jams) to control or disrupt the network, compromising safety by creating unreliable traffic info or forcing bad routes. Attackers forge IDs by stealing legitimate ones or creating new ones, aiming to manipulate decisions in these decentralized systems, requiring strong authentication and behavior analysis for mitigation.
  2. 2.
    Data Integrity and Freshness Attacks:
    • Position/Velocity Data Falsification Attack: where a malicious vehicle broadcasts false location/speed data within its basic safety message (BSM). This attack can mislead other vehicles, potentially leading to accidents or congestion by creating a false sense of their position.
    • Replay Attack: involves an attacker capturing legitimate vehicle data and retransmitting it later to deceive the system, potentially causing accidents.
    • Spoofing Attack: involves a malicious actor sending fake signals to trick the vehicle’s sensors, leading it to a false location or making it misinterpret its surroundings. Sometimes, this attack is preceded by an impersonation attack. Examples of such attack could be position/velocity data falsification attack, GPS spoofing, LiDAR spoofing or other sensors’ data spoofing.
    • Man-in-the Middle (MiTM) Attack: occurs when a malicious entity secretly intercepts, relays, and potentially alters communication between legitimate nodes, such as vehicles, or roadside units (RSUs). It can be passive, eavesdropping or active, in which the attacker actively intervenes in the communication by dropping, delaying, or tampering with the content of messages, i.e. spoofing attack.
    Figure 2: Autonomous Vehicles Zero-Trust Decentralized Blockchain-based System Example Target Attacks
  3. 3.
    Collusion and Coordinated Attacks:
    • Deception and Misinformation: Malicious vehicles or roadside units can collude to broadcast false information, such as forged warnings about obstacles or accidents, leading a victim vehicle to react incorrectly.
    • Sensor Spoofing and Jamming: Multiple vehicles can coordinate to spoof sensors, such as radar or GPS, simultaneously to confuse a target vehicle’s perception system or jam its signals, creating a more robust and difficult-to-detect attack.
    • Adversarial Machine Learning: Attackers can craft specific, coordinated inputs to confuse the machine learning models used for decision-making, such as altering perception data from multiple sensors to cause a misclassification.
  4. 4.
    Availability and Network-Level Attacks:
    • Distributed Denial-of-Service (DDOS) Attack: is a malicious attempt to disrupt the normal functionality and availability of blockchain network, vehicle, or RSU by overwhelming it with a flood of traffic from multiple coordinated sources.
  5. 5.
    Accountability and Trust Attacks:
    • Repudiation Attacks: happen when a system lacks strong controls to prove who initiated an action, allowing a malicious actor to deny sending a message or performing a transaction, or to alter logs to frame someone else, making digital records unreliable. The goal of such attacks could be denial of origin, log temparing or data forgery.
    • Single Point of Failure Attacks: target a single component, like a critical sensor, communication link, or central server, that, if compromised or disabled, brings down the entire system or causes catastrophic failure, such as a crash or loss of control. Attackers exploit these vulnerabilities through sensor spoofing, V2X (Vehicle-to-Everything) exploits, or denial-of-service (DOS) attacks.
These attacks have been documented in prior work on vehicular networks and cooperative driving including [yousseef2025autonomous, miller2015remote, amoozadeh2015security, satam2022autonomous], motivating zero-trust based decentralized and continuously verifiable identity management solutions. Figure 2 illustrates some of these target attacks. The proposed D-IM framework mitigates these attacks through decentralized verification, hash-based integrity checks, and continuous authentication.
4 DECENTRALIZED IDENTITY MANAGEMENT (D-IM) SYSTEM ARCHITECTURE
@@
An overview of the system steps is shown in Figure 3. The overall communication framework integrates Zero Trust principles into V2X security. Authentication via blockchain precedes all information sharing, ensuring that only legitimate participants contribute to cooperative V2X operations. Each vehicle and RSU runs a local copy of the framework code and communicates with the blockchain to achieve secure and verifiable V2X communications. The framework enforces Zero Trust by requiring identity verification for every interaction, eliminating reliance on any single trusted entity. This paper focus on the authentication and authorization of joining vehicles (Step 1). The trust consensus voting and smart contract-based penalization (steps 3 and 5) are reserved for future work.
Table 2: Algorithm 1 used Parameters
Parameter Explanation
$ID_{1}$ Vehicle 1 Pseudonymous Identity
$ID_{2}$ Vehicle 2 Pseudonymous Identity
TSTS Timestamp
LOCLOC Vehicle’s location
$K_{pb_{1}}$ Vehicle 1 Public Key
$K_{pvt_{1}}$ Vehicle 1 Private Key
$K_{pb_{2}}$ Vehicle 2 Public Key
$K_{pvt_{2}}$ Vehicle 2 Private Key
$Q$ , $R$ Random Numbers used to generate Diffie-Hellman based secret key
$K_{s}$ Diffie-Hellman Secret Key
@@
11:   Both vehicles compute $K_{s}$ from $Q$ and $R$ via Diffie-Hellman
@@
The detailed algorithm is provided in Algorithm 1 with its parameters summarized in Table 2. The algorithm formalizes the identification and authorization steps between communicating nodes. The protocol operates as follows:
@@
    Initialization: Each vehicle submits its pseudonymous ID, public key, timestamp (to ensure message freshness), location (to verify physical proximity to nearby RSUs to ensure that a vehicle is physically present at its claimed location using received signal strength (RSS) measurements and to enable reliable association of reported events with a concrete geographic context), and the hash of its ID with its private key to the blockchain. The hash of the vehicle’s ID and private key is used to verify the real identity of the vehicle since the public/private key pair is associated with real vehicle plate number or other identifying information.
    • Step 1: Vehicle1 submits $(ID_{1},K_{pb_{1}},TS,Loc,Hash(ID_{1}||K_{pvt_{1}}))$ to blockchain.
    • Step 2: Vehicle2 submits $(ID_{2},K_{pb_{2}},TS,Loc,Hash(ID_{2}||K_{pvt_{2}}))$ to blockchain.
@@
    Request/Response: When Vehicle 1 requests to communicate with Vehicle 2, it queries the blockchain for Vehicle 2’s credentials and gets the blockchain response with Vehicle 2’s credentials.
    • Step 3: Vehicle1 requests $ID_{2}$ info from blockchain.
    • Step 4: blockchain responds with $(ID_{2},K_{pb_{2}},TS,Loc,Hash(ID_{2}||K_{pvt_{2}}))$
    Then, Vehicle 1 initiates communication with Vehicle 2 sending its own ID $ID_{1}$ to announce who is initiating communication, Vehicle 2’s ID $ID_{2}$ for Vehicle 2 to know it is the intended recipient, its own public key $K_{pb_{1}}$ to be used for future communication, current timestamp $TS$, random number $Q$ to be used to generate Diffie-Hellman’s secret key, and the hash $Hash(ID_{2}||K_{pvt_{2}}$ to verify Vehicle 1 is who claims it is since only the blockchain knows this hash and hence Vehicle 1 one must have be authorized by the blockchain to get the hash. The whole message is encrypted using Vehicle 2’s public key $K_{pb_{2}}$.
[…]
    Similarly, Vehicle 2 queries the blockchain for Vehicle 1’s information and gets the blockchain response with Vehicle 2’s credentials.
    • Step 6: Vehicle2 requests $ID_{1}$ info from blockchain
    • Step 7: blockchain responds with $(ID_{1},K_{pb_{1}},TS,Loc,Hash(ID_{1}||K_{pvt_{1}}))$
@@
    Verification: Vehicle 2 cross-verifies the Vehicle 1’s public key against blockchain records and received hash against its local records to ensure authenticity.
[…]
    If the received key and hash match, Vehicle 2 sends to Vehicle 1 its own ID $ID_{2}$ to announce who is initiating communication, Vehicle 1’s ID $ID_{1}$ for Vehicle 1 to know it is the intended recipient, its own public key $K_{pb_{2}}$, current timestamp $TS$, random number $R$ to be used to generate Diffie-Hellman’s secret key, and the hash $Hash(ID_{2}||Q))$ to be used to verify Vehicle 2 received the right $Q$ value. The whole message is encrypted using Vehicle 1’s public key $K_{pb_{1}}$.
[…]
    Vehicle 1 cross-verifies received hash against its local computed hash to ensure authenticity.
    • Step 10: Hash($ID_{2}$||Q) received = locally computed Hash($ID_{2}$||Q)
@@
    Session Key Establishment: Both vehicles have exchanged random values ($Q$, $R$) encrypted with the other’s public key, enabling them to compute a shared session key $K_{s}$ using the Diffie-Hellman method.
    • Step 11: Both vehicles compute $K_{s}$ from $Q$ and $R$ via Diffie-Hellman